OpenStudy (anonymous):

Find the exact area of the circle with the given radius. r = 11.64 m A. 11.64π m² B. 23.28π m² C. 33.8724π m² D. 135.4896π m²

OpenStudy (anonymous):

A = pi*r^2

OpenStudy (anonymous):

\[11.64^2 = ?\]

OpenStudy (anonymous):

oh I got D

OpenStudy (anonymous):

Yeah that's right.

OpenStudy (anonymous):

A? or D?

OpenStudy (anonymous):

I need to know the exact right answer this thing is stupid! :P

OpenStudy (anonymous):

\[A = \pi r^2\] area of a circle.

OpenStudy (anonymous):

So just plug in for r, and square it you'll get D.

OpenStudy (anonymous):

THX!!!!!!

OpenStudy (anonymous):

THEORETICALLY YOU BOTH GET A MEDAL!

OpenStudy (anonymous):

No Problem, Thanks, Hope It Helped! =)

OpenStudy (anonymous):

Btw @angelbaby_2434 it's not looking for the equation of a circle, it's looking for the area.

OpenStudy (anonymous):

Oh okay sorry =( Not so great at math sorry, was jus trying to help but sorry if I messed anyone up, I didn't mean to! sorry!

OpenStudy (anonymous):

No worries :) was just trying to clarify the problem for you

OpenStudy (anonymous):

Thank You! =) it helps, so I know next time =)